OpenStudy (anonymous):

The work of a student to solve the equation 2(3x − 4) = 8 + 2x + 4 is shown below: Step 1: 2(3x − 4) = 8 + 2x + 4 Step 2: 5x − 6 = 12 + 2x Step 3: 5x − 2x = 12 + 6 Step 4: 3x = 18 Step 5: x = 6 In which step did the student first make an error and what is the correct step? Step 2; 6x − 6 = 2(6 + x + 2) Step 2; 6x − 8 = 12 + 2x Step 3; 5x − 2x = 12 − 6 Step 3; 5x + 2x = 12 + 6

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

Step 1: 2(3x − 4) = 8 + 2x + 4 Step 2: 5x − 6 = 12 + 2x Step 3: 5x − 2x = 12 + 6 Step 4: 3x = 18 Step 5: x = 6

OpenStudy (anonymous):

The work of a student to solve the equation 2(3x − 4) = 8 + 2x + 4 is shown below: Step 1: 2(3x − 4) = 8 + 2x + 4 Step 2: 5x − 6 = 12 + 2x Step 3: 5x − 2x = 12 + 6 Step 4: 3x = 18 Step 5: x = 6 In which step did the student first make an error and what is the correct step? Step 2; 6x − 6 = 2(6 + x + 2) Step 2; 6x − 8 = 12 + 2x Step 3; 5x − 2x = 12 − 6 Step 3; 5x + 2x = 12 + 6

OpenStudy (lynfran):

2(3x-4)=8+2x+4 6x-8=12+2x 6x-2x=12+8 4x=20 x=20/4 x=5 this is how this question shd be worked...

OpenStudy (anonymous):

thanks can you help me with another

OpenStudy (lynfran):

sure

OpenStudy (anonymous):

Which of these is a simplified form of the equation 6y + 4 = 8 + 2y + 2y? 6y = 8 10y = 12 2y = 4 4y = 12

OpenStudy (lynfran):

6y+4=8+2y+2y 6y+4=8+4y 6y-4y=8-4 2y=4 y=4/2 y=2

OpenStudy (lynfran):

can u see it??

OpenStudy (lynfran):

check the working i just did and look for the answer

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

thanksss can u help with more

OpenStudy (lynfran):

sure

OpenStudy (anonymous):

How many solutions does the equation −3a + 3a + 6 = 7 have? One Two Infinitely many None

OpenStudy (lynfran):

-3a+3a=0 so none

OpenStudy (anonymous):

What is the solution to the equation 1.6m − 4.8 = −1.6m? m = 0.5 m = 0.7 m = 1.5 m = 3

OpenStudy (lynfran):

the a term cancel out each other

OpenStudy (lynfran):

1.6m-4.8=-1.6m 1.6m+1.6m=4.8 3.2m=4.8 m=4.8/3.2 m=1.5

OpenStudy (anonymous):

wowww ur amazing

OpenStudy (anonymous):

would i be pushing it if i ask for more help?

OpenStudy (lynfran):

nah

OpenStudy (anonymous):

What is the value of x in the equation −6 + x = −3? −9 −3 3 9

OpenStudy (anonymous):

aye u got insta?

OpenStudy (lynfran):

-6+x=-3 x=-3+6 x=3

OpenStudy (anonymous):

What is the value of x in the equation −x = 4 − 3x + 6? 5 10 −5 −10

OpenStudy (anonymous):

Which statement is true about the equation fraction 3 over 4z = fraction 1 over 4z − 3 + 5? It has no solution. It has one solution It has two solutions. It has infinitely many solutions.

OpenStudy (lynfran):

-x=4-3x+6 -x=10-3x -x+3x=10 2x=10 x=5

OpenStudy (anonymous):

so D?

OpenStudy (lynfran):

What is the value of x in the equation −x = 4 − 3x + 6? x=5

OpenStudy (lynfran):

\[\frac{ 3 }{ 4z }=\frac{ 1 }{ 4z }-3+5 \] is this the correct question/equation

OpenStudy (anonymous):

How many solutions does the equation 6a − 3a − 6 = −2 + 3 have? Two None One Infinitely many

OpenStudy (lynfran):

im talking about this question we havent started as yet What is the value of x in the equation −x = 4 − 3x + 6? It has no solution. It has one solution It has two solutions. It has infinitely many solutions.

OpenStudy (lynfran):

so this was the question Which statement is true about the equation fraction 3 over 4z = fraction 1 over 4z − 3 + 5?

OpenStudy (lynfran):

we havent started the above question as yet

OpenStudy (lynfran):

but for this question How many solutions does the equation 6a − 3a − 6 = −2 + 3 have? the answer is one

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

What is the value of z in the equation 2(4z − 3 − 1) = 166 − 46? 14 15 16 17

OpenStudy (lynfran):

2(4z-3-1)=166-46 8z-6-2=120 8z=120+8 z=16

OpenStudy (anonymous):

hey can u help with more

OpenStudy (lynfran):

ok 3more but you have to look at the pattern of my solution when it comes to solving equation ok

OpenStudy (anonymous):

k thxs

OpenStudy (anonymous):

How many solutions does the equation −3y + 3y + 4 = 4 have? Zero Infinitely many One Four

OpenStudy (lynfran):

-3y+3y+4=4 and -3y+3y=0 so no solution

OpenStudy (vheah):

The student committed an error on Step 2 because instead of multiplying 2 to 3x the student ended up adding the two, getting 5x. Correct step is 6x-8 = 12x + 2x